QUESTION 17 The following statistics were calculated Treatment 2 14 35 Statistic /2 10 30 10 40 10 10 10 In the calculations to complete an ANOVA table it was found that In the further calculations given below, find the SST 737.9 and SSE 490.. one incorrect statement 1. The numbe 2. The 3. MST 264.0 4. MSE 10.0 5. F 24.60 r of degrees of freedom in the mean square for treatments is 3 number of degrees of freedom in the mean square for error is 49.
Answer #1

Question 17)

Given that,

Sum of square due to treatment (SST) = 737.9

Sum of square due to error (SSE) = 490.0

Total sample size (n) = 10 + 14 + 11 + 18 = 53

Therefore, degrees of freedom for treatment = 4 -1 = 3

(since there are 4 treatments)

Total degrees of freedom = n - 1 = 53 - 1 = 52

Error degrees of freedom = Total df - Treatment df = 52 - 3 = 49

Now,

MST = SST/Treatment df = 737.9/3 = 246.0

MSE = SSE/Error df = 490.0/49 = 10.0

and

F = MST/MSE = 246.0/10.0 = 24.60

Hence incorrect statement is,

3) MST = 264.0